OpenStudy (anonymous):

use factoring to solve this equation x^2-12x+27=0

OpenStudy (anonymous):

x^2-12x+27=0 x^2 -9x -3x + 27 = 0 x(x-9) -3(X-9) = 0 (x-3)(x-9) = 0 This is only possible if either x-3=0 or x-9 = 0 if x-3 = 0 ==> x = 3 if x-9 = 0 ==> x = 9 so solution for the given quadratic equation is x = 3 , 9